x=2y+z=2
5x=z
y=-2x

Answer by mangopeeler07(462)

x=2y+z=2
5x=z
y=-2x

Plug in y and z
x=2(-2x)+5x=2

x=-4x+5x=2
x=x=2

x=2

5(2)=z
z=10

y=-2(2)
y=-4

Answer:
x=2
y=-4
z=10